OBSERVATION OF COMPETING PATHWAYS IN THE RELAXATION OF ICl* IN A He SUPERSONIC EXPANSION

JOSHUA P. DARR and \underline{RICHARD A. LOOMIS}, Department of Chemistry, Washington University, One Brookings Drive, CB 1134, Saint Louis, MO 63130.

Collision-induced relaxation of ICl* and differing caging mechanisms of electronically excited ICl* within He⋯ICl* complexes are investigated using laser-induced fluorescence and two-laser, pump-probe spectroscopy. It is found that collisions of He atoms with molecules prepared in the short-lived ICl(B ^{3}Π_{0+},v′=3) rotor states occurs on timescales sufficient for vibrational relaxation down to the long-lived ICl(B,v′=2) rotor states. It is also shown that when linear He⋯ICl(X ^{1}Σ_{0+},v′′=0) complexes are promoted to the continuum of states correlating with the repulsive, inner wall of the ICl(A ^{3}Π_{1}) potential, the dissociating ICl(A) molecule undergoes kinematic one-atom caging, ejecting the weakly bound He atom, and re-stabilizing the molecule in high lying vibrational levels within the A electronic state with very little rotational excitation. Features associated with transitions of the linear He⋯I^{35}Cl(X,v′′=0) complex to intermolecular states associated with the ICl(B) diabat that are coupled with those in the ICl(B′) adiabat formed by an avoided curve crossing are also observed in laser-induced fluorescence spectra. The ICl* molecules within the excited state complex undergoes a non-adiabatic one-atom caging mechanism forming I^{35}Cl(B,v′) products with little rotational excitation.